OpenStudy (anonymous):

-2x(3X^2+X-5) MULTIPLYING @-------College Al…

OpenStudy (anonymous):

Mind rewriting the problem more legibly?

OpenStudy (anonymous):

-2x(3x^2+x-5)

OpenStudy (anonymous):

-6x^3-2x^2+10x

OpenStudy (anonymous):

-2x * 3x^2 = -6x^3 -2x * x = -2x^2 -2x * (-5) = 10x Put them together including their signs and you have -6x^3 - 2x^2 + 10x